(ii) 2x – 1/6 = 3
(ii) 2x – 1/6 = 3
(ii) 2x – 1/6 = 3
(ii) 2x – 1/6 = 3
(ii) 2x – 1/6 = 3
(ii) 2x – 1/6 = 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x – 1/6 = 3

2x = 3 + 1/6

2x = 19/6

x = 19/12

Thanks